OpenStudy (anonymous):

A year ago we placed some money in a bank account with an annual simple interest of 13%. How much did we deposit there a year ago if now there is $1356 in the account?

OpenStudy (anonymous):

I'm lost as to how to approach the problem, can someone please help?

OpenStudy (tkhunny):

I = Prt Always think about this. This is what Simple Interest is all about. We have r = 0.13 I = P(0.13)t We have t = 1 I = P(0.13)(1) = P(0.13) We have 1356. This is made up of the original deposit, P, and the Interest Earned, I 1356 = P + I = P + P(0.13) = P(1.13) One step at a time. Can you finish?

OpenStudy (anonymous):

We're its 1200 thanks! I got it now!